OpenStudy (anonymous):

In a regular pentagon ABCDE and regular hexagon MNOPQR, 2AB = 3MN and CD + QR = 10. Which is 4BC + 5PQ? Answer 44 42 38 33

OpenStudy (anonymous):

2 ab = 3 mn ab + mn = 10 ( as ab = cd and mn = qr) ab = 3mn / 2 3mn / 2 + mn = 10 3mn + 2mn = 20 mn = 4 so ab = 6 4bc + 5pq = 24 + 20 = 44

OpenStudy (anonymous):

because ab = bc and mn = pq